The difference between the length and breadth of a rectangle is 12m. The perimeter of the rectangle is 144m. Find its dimensions?

Answer by checkley79(3341)

L=W-12
P=2L+2W
144=2(W-12)+2W
144=2W-24+2W
144=24=4W
168=4W
W=168/4
W=42 ANS.
L=42-12=30 ANS.
PROOF:
144=2*30+2*42
144=60+84
144=144